# Stochastic Multi-round Submodular Optimization with Budget

### v1

@@ 2.2 Maximization via Adaptive Policies
An interesting class of the above policies is that of semi-adaptive policies, in which the number of seeds that must be selected at each round (i.e., the budget per round) is determined non-adaptively in advance, and then the adaptive selection of items is performed within each round independently. Semi-adaptive policies are computationally simpler than general adaptive ones, but it might have worse performance guarantees. Actually, we will show that this worsening is limited for our problem.
@@
2.3 Properties of the State Functions
In the following, we list some additional properties that the state functions must verify in our setting.
The first subset of properties will focus on the complexity of the state function. Specifically, we assume that each state function $f_{t}$ is representable and computable in polynomial time. Moreover, we also constraint the maximum value of a state function, as follows.
@@ 2.3 Properties of the State Functions
A state function $f_{t}$ is polynomially bounded if there exists real values $\lambda,\Lambda>0$, where $\Lambda$ and $1/\lambda$ are polynomially bounded in the size of $I$, such that the following conditions hold: (ii) $f_{t}(\{v\}\cup S_{t},\eta_{t})-f_{t}(S_{t},\eta_{t})\leq\Lambda$ for any $v\in V$, $S_{t}\subseteq V$ and state $\eta_{t}$; (ii) there exist $t\in[T]$ and $v\in V$ such that $\lambda\leq\mathbb{E}_{\bm{\eta}_{t}\mathcal{P}_{t}}[f_{t}(\{v\},\bm{\eta}_{t})]$.
@@
We further assume that state functions satisfy a certain property connected with the notion of submodularity. Given two partial states $\langle S_{t}\rangle:=\langle S_{t},\eta_{t}\rangle$ and $\langle S^{\prime}_{t}\rangle:=\langle S^{\prime}_{t},\eta^{\prime}_{t}\rangle$, we say that $\langle S_{t}\rangle$ is a sub-state of $\langle S^{\prime}_{t}\rangle$ (equivalently, we say $\langle S_{t}\rangle\preceq\langle S^{\prime}_{t}\rangle$) if $S_{t}\subseteq S^{\prime}_{t}$, and $\eta_{t}(e)=\eta^{\prime}_{t}(e)$ for any $e\in S_{t}$, that is, the local states observed in $\langle S_{t}\rangle$ can be also observed in $\langle S^{\prime}_{t}\rangle$. Given a partial state $\langle S_{t}\rangle$ at round $t$ and $v\in V$, let $\Delta_{t}(v\mid\langle S_{t}\rangle):=\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}}\left[f_{t}(S_{t}\cup\{v\},\bm{\eta}_{t})-f(S_{t},\bm{\eta}_{t})|\langle S_{t}\rangle\preceq\bm{\eta}_{t}\right]$ i.e., the expected increment of $f_{t}$ obtained when adding item $v$ to $S_{t}$, conditioned by the observation of partial state $\langle S_{t}\rangle$.
@@
A state function $f_{t}:2^{V}\rightarrow{\mathbb{R}}_{\geq 0}$ is adaptive submodular if, for all $\langle S_{t}\rangle\preceq\langle S^{\prime}_{t}\rangle$ and $v\in V$ we get $\Delta_{t}(v\mid\langle S_{t}\rangle)\geq\Delta_{t}(v\mid\langle S^{\prime}_{t}\rangle)$.
Henceforth, we assume that state functions satisfy all the properties described above.
@@ 2.4 The Stochastic Budgeted Multi-round Submodular Maximization Problem
In particular, we will focus on policies with guaranteed approximation, defined as follows. Given an instance $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in[T]})$ of SBMSm, $\alpha\in[0,1]$, and a multi-round policy $\pi$ for $I$, we say that $\pi$ is $\alpha$-approximate, if $\sigma(\pi)\geq\alpha\cdot\sigma(\pi^{*})-\beta$, where $\pi^{*}$ denotes an optimal multi-round policy for $I$.
In the remainder of the paper, we will implicitly assume that $B<nT$, where $B$ is the budget of the input instance. Indeed, if $B\geq nT$, we could select a set $V$ at each round so that the state function is trivially maximized at each round. Furthermore, the assumption $B<nT$ implies that the budget $B$ is polynomial in the size of the input instance.
@@ 2.5.1 Influence Maximization
The Independent Cascade diffusion model (IC) [26], given a graph $G=(V,E)$, a probability $p\colon E\rightarrow[0,1]$ defined on edges, an obstruction $F\subseteq V$, and a set of seeds $S_{0}\subseteq V\setminus F$, returns a set $A$ of active nodes computed as follows: for each round $t\geq 0$ such that $S_{t}\neq\emptyset$, $S_{t}$ is added to $A$, and then for each edge $(u,v)\in E$ such that $u\in S_{t}$ and $v\notin A\cup F$, add $v$ to $S_{t+1}$ with probability $p(u,v)$.
@@
In the problem of Influence Maximization under Independent Cascade diffusion model (IM-IC) [26] we are given a graph $G=(V,E)$, a weight function $w\colon V\rightarrow\mathbb{R}_{+}$ defined on vertices, a probability $p\colon E\rightarrow[0,1]$ defined on edges, and a budget $B$, and we are asked to select the set $S^{*}$ of at most $B$ vertices that maximizes the expected sum of weights of active nodes returned by the Independent Cascade diffusion model on input $G$, $p$, an empty obstruction and $S$, i.e., $S^{*}=\arg\max_{S\subseteq V\colon|S|\leq B}\mathbb{E}_{p}\left[\sum_{v\in IC(G,p,\emptyset,S)}w(v)\right]$.
@@
Recently, an adaptive variant of the problem has been proposed where seeds are introduced one at a time, using an adaptive policy. To define the adaptive variant of the problem we need first to define the concept of adaptive policy. Given a graph $G=(V,E)$, a probability $p\colon E\rightarrow[0,1]$ defined on edges, and a budget $B$, an adaptive policy $\pi$ for IM-IC works as follows: it selects a first vertex $v_{1}$, observes the set $A_{1}$ of vertices activated by seed $v$, i.e., $A_{1}=IC(G,p,\emptyset,\{v_{1}\})$, then selects a new vertex $v_{2}\in V\setminus A_{1}$, observes the set $A_{2}$ of newly activated vertices, i.e., $A_{2}=IC(G,p,A_{1},\{v_{2}\})$, then selects a new vertex $v_{3}\in V\setminus(A_{1}\cup A_{2})$, observes the set $A_{3}$ of newly activated vertices, i.e., $A_{3}=IC(G,p,A_{1}\cup A_{2},\{v_{3}\})$, and so on until either there is no vertex that can be selected or $B$ vertices have been already selected. The set of nodes activated through an adaptive policy $\pi$ is $A(\pi)=\bigcup_{i\in[B]}A_{i}$.
@@
In the problem of Adaptive Influence Maximization under Independent Cascade diffusion model [22] we are given a graph $G=(V,E)$, a weight function $w\colon V\rightarrow\mathbb{R}_{+}$ defined on vertices, a probability $p\colon E\rightarrow[0,1]$ defined on edges, and a budget $B$, and we are asked to select the adaptive policy $\pi^{*}$ that maximizes the expected sum of weights of activated nodes, i.e., $\pi^{*}=\arg\max_{\pi}\mathbb{E}_{p}\left[\sum_{v\in A(\pi)}w(v)\right]$.
@@
In this paper we are interested to multi-round settings. Hence, we need to extend previous concepts to the multi-round setting. Given a graph $G=(V,E)$, a probability $p_{t}\colon E\rightarrow[0,1]$ defined on edges (for each round $t$), a budget $B$, and a time horizon $T$, a multi-round adaptive policy $\pi$ for IM-IC works as follows: at each round $t\in[T]$, let $S^{t}(\pi)$ be the set of seeds selected by $\pi$ in round $t$ and $A^{t}_{v}(\pi)$ be the set of nodes activated by $v\in S^{t}(\pi)$ (we will next omit the reference to $\pi$ when clear from the context). Note that both $S^{t}$ and $A^{t}_{v}$ are initially empty. Then, the policy at each round $t\in[T]$, if the total number of selected vertices is lower than $B$,
@@
    decides whether to choose another seed in the current round or to move to the next round (note that if all vertices have been activated in the current round, then the policy must necessarily choose to move to the next round);
@@
    in the first case, it selects a vertex $v$ that has not been yet activated in the current round, i.e., $v\notin A^{t}_{S^{t}}=\bigcup_{u\in S^{t}}A^{t}_{u}$, observes the set $A^{t}_{v}$ of newly activated vertices, i.e., $A^{t}_{v}=IC(G,p,A^{t}_{S^{t}},\{v\})$, adds $v$ to $S^{t}$, and repeat from step 1.
Finally, in the problem of Multi-Round Adaptive Influence Maximization under Independent Cascade diffusion model (MR-IM-IC) we are given a graph $G=(V,E)$, a time horizon $T$, a weight function $w_{t}\colon V\rightarrow\mathbb{R}_{+}$ defined on vertices for every $t\in[T]$, a probability $p_{t}\colon E\rightarrow[0,1]$ defined on edges for each round $t\in[T]$, and a budget $B$, and we are asked to select the multi-round adaptive policy $\pi^{*}$ that maximizes the expected sum of weights of activated nodes over $t$ rounds, i.e., $\pi^{*}=\arg\max_{\pi}\varsigma(\pi)$ where $\varsigma(\pi)=\mathbb{E}_{(p_{t})_{t\in[T]}}\left[\sum_{t\in[T]}\sum_{v\in A^{t}_{S^{t}}(\pi)}w_{t}(v)\right]$.
We stress that our definition allows that both vertices’ weights and edge probabilities may change from one round to the next: this can be useful to implement a variable discount factor (e.g., we care more about the number of people influenced in the last phases of the campaign) or to implement a dynamic environment in which the members of a population are known to appear and disappear over time, or the relevance of population members changes over time (e.g., in liquid democracy settings [6] at the beginning we are more interested in influencing a large portion of the population for manipulating the choice of delegates, while later in the campaign we are more concerned in influencing those members that have more chances to be chosen as delegates, in order to manipulate the outcome of their voting process).
@@
We start by showing how an instance of the MR-IM-IC problem can be mapped to an instance of SBSMm. Let $M=(T,B,V,E,(p_{t})_{t\in[T]},(w_{t})_{t\in[T]})$ be an instance of MR-IM-IC and let $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in[T]})$ be an instance of SBSMm. The mapping works as follows:
  • $T$ and $B$ are exactly the same as in $M$ (hence, we use exactly the same name);
  • the set $V$ of items is the set of nodes in the social network;
@@
  • the set $H(v)$ of possible local states of $v$ corresponds to the possible sets of nodes influenced by an independent cascade starting from node $v$, i.e., $H(v)=\{X\subseteq V\mid v\in X\}$;
  • the probability $\mathcal{P}_{t}(\eta_{t})$ is the probability that an independent cascade influences exactly the $\eta_{t}(v)\in H(v)$ nodes when started from $v$ for every $v\in V$. Specifically, we can define $\mathcal{P}_{t}$ as follows: for each spanning subgraph $G_{E^{\prime}}$ of $G$, i.e., $G_{L}=(V,E^{\prime})$ with $E^{\prime}\subseteq E$, we set $P_{t}(G_{E^{\prime}})=\prod_{e\in E\cap E^{\prime}}p_{t}(e)\prod_{e\in E\setminus E^{\prime}}(1-p_{t}(e))$, i.e. $G_{E^{\prime}}$ denotes one of the possible live graphs and $P(G_{E^{\prime}})$ its probability; for each $v$ and each $E^{\prime}\subseteq E$, we denote as $R_{E^{\prime}}(v)$ the set of nodes reachable from $v$ in $G_{E^{\prime}}$; for a state $\eta_{t}$, we set $\Xi(\eta_{t})=\{E^{\prime}\subseteq E\mid\eta_{t}(v)=R_{E}^{\prime}(v)\forall v\in V\}$, i.e., $\Xi(\eta_{t})$ is the set of live graph that are compatible with the state of all vertices described in $\eta_{t}$; we say that $\eta_{t}$ is feasible if $|\Xi(\eta_{t})|\geq 1$; then, $\mathcal{P}_{t}(\eta_{t})=0$ if $\eta_{t}$ is infeasible, and $\mathcal{P}_{t}(\eta_{t})=\sum_{E^{\prime}\in\Xi(\eta_{t})}P_{t}(G_{E^{\prime}})$;
  • the state function $f_{t}$ is defined as $f_{t}(S,\eta_{t})=\sum_{v\in\bigcup_{s\in S}\eta_{t}(s)}w_{t}(v)$.
@@
According to this mapping between $M$ and $I$, and the obvious mapping between a multi-round policy $\pi$ for IM-IC and a multi-round policy $\pi^{\prime}$ for SBMSm, it is immediate to check that $\varsigma(\pi)=\sigma(\pi^{\prime})$, and thus finding the policy maximizing the latter is equivalent to find a policy maximizing the former.
@@
In order to establish that $I$ is a valid instance for SBMSm we need also to prove that $f_{t}$ is polynomially computable and representable, it is polynomially bounded, and it satisfies monotonicity and adaptive submodularity.
Polynomial representation and computation immediately follows from the definition above. As for polynomial boundedness, this property is immediately satisfied if there exist constants $c,C>0$ such that $w_{t}(v)\leq C$ for every $v$ and every $t$, and there exist a round $t$ and a node $v$ such that $c\leq w_{t}(v)$. In this case we have $\Lambda\leq Cn$ and $\lambda=c$.
Monotonicity is immediately satisfied since the information diffusion cannot decrease if the set of observed edges expands. As for the adaptive submodularity, it is not hard to check that this condition is satisfied: intuitively, for each realization of the live graph, as described above, the increment of $f_{t}$ due to $v$ is given by nodes reachable from $v$ in this live graph, but not reachable by previously selected seeds; and it is clearly impossible to expand this subset with more previously selected seeds.

### v2

@@ 2.2 Maximization via Adaptive Policies
An interesting class of the above policies is that of partially adaptive policies, in which the number of items that must be selected at each round (i.e., the budget per round) is determined non-adaptively in advance, and then the adaptive selection of items is performed within each round independently. Partially adaptive policies are computationally simpler than general adaptive ones, but it might have worse performance guarantees. Actually, we will show that this worsening is limited for our problem.
@@
2.3 Properties of the Objective Functions
In the following, we list some additional properties that the objective functions must verify in our setting.
The first subset of properties will focus on the complexity of the objective function. Specifically, we assume that each objective function $f_{t}$ is representable and computable in polynomial time. Moreover, we also constraint the maximum value of an objective function, as follows.
@@ 2.3 Properties of the Objective Functions
An objective function $f_{t}$ is polynomially bounded if there exist real values $\lambda,\Lambda>0$, with $\Lambda$ and $1/\lambda$ being polynomially bounded in the size of $I$, such that the following conditions hold: (i) $f_{t}(\{v\}\cup S_{t},\eta_{t})-f_{t}(S_{t},\eta_{t})\leq\Lambda$ for any $v\in V$, $S_{t}\subseteq V$ and state $\eta_{t}$; (ii) there exist $t\in[T]$ and $v\in V$ such that $\lambda\leq\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\{v\},\bm{\eta}_{t})]$.
@@
We further assume that objective functions satisfy a certain property connected with the notion of submodularity. Given two partial states $\langle S_{t}\rangle:=\langle S_{t},\eta_{t}\rangle$ and $\langle S^{\prime}_{t}\rangle:=\langle S^{\prime}_{t},\eta^{\prime}_{t}\rangle$, we say that $\langle S_{t}\rangle$ is a sub-state of $\langle S^{\prime}_{t}\rangle$ (equivalently, we say $\langle S_{t}\rangle\preceq\langle S^{\prime}_{t}\rangle$) if $S_{t}\subseteq S^{\prime}_{t}$, and $\eta_{t}(v)=\eta^{\prime}_{t}(v)$ for any $v\in S_{t}$, that is, the local states observed in $\langle S_{t}\rangle$ coincide with those observed in the restriction of $\langle S^{\prime}_{t}\rangle$ to set $S_{t}$; we observe that $\langle V,\eta_{t}\rangle=\eta_{t}$ holds for any state $\eta_{t}$ (by definition). Given a partial state $\langle S_{t}\rangle$ at round $t$ and $v\in V$, let $\Delta_{t}(v\mid\langle S_{t}\rangle):=\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}\left[f_{t}(S_{t}\cup\{v\},\bm{\eta}_{t})-f(S_{t},\bm{\eta}_{t})|\langle S_{t}\rangle\preceq\bm{\eta}_{t}\right]$ i.e., the expected increment of $f_{t}$ obtained when adding item $v$ to $S_{t}$, conditioned by the observation of partial state $\langle S_{t}\rangle$.
@@
An objective function $f_{t}:2^{V}\rightarrow{\mathbb{R}}_{\geq 0}$ is adaptive submodular if, for all $\langle S_{t}\rangle\preceq\langle S^{\prime}_{t}\rangle$ and $v\in V$ we get $\Delta_{t}(v\mid\langle S_{t}\rangle)\geq\Delta_{t}(v\mid\langle S^{\prime}_{t}\rangle)$.
Henceforth, we assume that objective functions satisfy all the properties described above.
@@ 2.4 The Stochastic Budgeted Multi-round Submodular Maximization Problem
In particular, we will focus on policies with guaranteed approximation, defined as follows. Given an instance $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in[T]})$ of SBMSm, $\alpha\in[0,1]$, and a multi-round policy $\pi$ for $I$, we say that $\pi$ is $\alpha$-approximate (or, equivalently, $\alpha$ is the approximation factor of $\pi$), if $\sigma(\pi)\geq\alpha\cdot\sigma(\pi^{*})$, where $\pi^{*}$ denotes an optimal multi-round policy for $I$;
We observe that the size of the input instance $I$ generally depends on the succinct representation of all the parameters defining $I$. For the sake of simplicity, in the remainder of the paper we will implicitly assume that the size of the input instance is polynomial in $n$ and $T$. The polynomial dependence in the number of items $n$ is a standard assumption/property of problems connected with the optimization of functions defined on subset of items. The polynomial dependence in $T$ holds since the input instance $I$ can be seen as a collection of $T$ single-round sub-instances (one for each round $t\in[T]$), which are assumed to be explicitly listed. Furthermore, we will implicitly assume that $B<nT$, where $B$ is the budget of the input instance. Indeed, if $B\geq nT$, we could select a set $V$ at each round so that the objective function is trivially maximized at each round. Thus, the assumption $B<nT$ implies that the budget $B$ is also polynomial in the size of $I$. Finally, we allow the size of the remaining parameters defining $I$ (i.e., local and global states, probability distributions and objective functions) to be exponential when explicitly represented, although the size of their succinct representation is not; this is the case, for instance, of influence maximization problems, which are described below.
Compatibly with several stochastic optimization problems, we also assume the existence of a polynomial-time oracle that works as follows: (i) for each $t\in[T]$, $S\subseteq V$ and state $\eta\in H$ given in input, it is able to compute $f_{t}(S,\eta)$; (ii) for each $t\in[T]$, sets $S,S^{\prime}$ with $S\subseteq S^{\prime}\subseteq V$ and partial state $\langle S\rangle$, the oracle is able to pick a partial state $\langle S^{\prime}\rangle\succeq\langle S\rangle$, drawn from the distribution $\mathcal{P}_{t}$ conditioned by the observation of $\langle S\rangle$.
@@ 2.5.1 Influence Maximization
The Independent Cascade diffusion model (IC) [32], given a graph $G=(V,E)$, a probability $p\colon E\rightarrow[0,1]$ defined on edges and a set of seeds $S_{0}\subseteq V$, returns a set $A=IC(G,p,S_{0})$ of active nodes iteratively computed over multiple (discrete) diffusion rounds [^2] Here, “diffusion round” has a different meaning from the term “round” considered in SBMSm. as follows: (i) at diffusion round 00, the set of activated nodes $A_{0}$ coincides with $S_{0}$; (ii) given a diffusion round $h\geq 1$, the set of nodes $A_{h}$ activated at time step $h$ is determined by including in $A_{h}$, for each edge $(u,v)$ with $u\in A_{h-1}$, the node $v$ with probability $p_{(u,v)}$, independently of the other edges; (iii) when the diffusion process reaches a diffusion round $k$ with $A_{k}=\emptyset$, that is, no further node can be activated, the diffusion process ends and the overall set of active nodes is $A:=\bigcup_{h=0}^{k-1}A_{h}$. We observe that, in the IC model, each active node has only one chance to activate each of its neighbors, the activation process associated with each edge is independent from the others.
@@
In the problem of Influence Maximization under Independent Cascade diffusion model (IM-IC) [32] we are given a graph $G=(V,E)$, a weight function $w\colon V\rightarrow\mathbb{R}_{\geq 0}$ defined on nodes, a probability $p\colon E\rightarrow[0,1]$ defined on edges, and a budget $B$, and we are asked to select the set $S^{*}$ of at most $B$ nodes that maximizes the expected sum of weights of active nodes returned by the IC diffusion model on input $G$ and $p$ i.e., $S^{*}=\arg\max_{S\subseteq V\colon|S|\leq B}\mathbb{E}_{p}\left[\sum_{v\in IC(G,p,S)}w(v)\right]$.
@@
Recently, an adaptive variant of the problem has been proposed where seeds are introduced one at a time, using an adaptive policy. To define the adaptive variant of the problem we need first to define the concept of adaptive policy. For any graph $G=(V,E)$ and set $U\subseteq V$ of nodes, let $G[U]$ denote the subgraph of $G$ induced by the nodes in $U$. Given a graph $G=(V,E)$, a probability $p\colon E\rightarrow[0,1]$ defined on edges, and a budget $B$, an adaptive policy $\pi$ for IM-IC works as follows: it selects a first vertex $v_{1}$, observes the set $A_{1}$ of nodes activated by seed $v$, i.e., $A_{1}=IC(G,p,\{v_{1}\})$, then selects a new vertex $v_{2}\in V\setminus A_{1}$, observes the set $A_{2}$ of newly activated nodes, i.e., $A_{2}=IC(G[V\setminus A_{1}],p,\{v_{2}\})$, then selects a new node $v_{3}\in V\setminus(A_{1}\cup A_{2})$, observes the set $A_{3}$ of newly activated nodes, i.e., $A_{3}=IC(G[V\setminus(A_{1}\cup A_{2})],p,\{v_{3}\})$, and so on until either there is no node that can be selected or $B$ nodes have been already selected. The set of nodes activated through an adaptive policy $\pi$ is $A(\pi)=\bigcup_{i\in[B]}A_{i}$.
@@
In the problem of Adaptive Influence Maximization under Independent Cascade diffusion model [26] we are given a graph $G=(V,E)$, a weight function $w\colon V\rightarrow\mathbb{R}_{+}$ defined on nodes, a probability $p\colon E\rightarrow[0,1]$ defined on edges, and a budget $B$, and we are asked to select the adaptive policy $\pi^{*}$ that maximizes the expected sum of weights of activated nodes, i.e., $\pi^{*}=\arg\max_{\pi}\mathbb{E}_{p}\left[\sum_{v\in A(\pi)}w(v)\right]$.
@@
In this paper we are interested to multi-round settings. Hence, we need to extend previous concepts to the multi-round setting. Given a graph $G=(V,E)$, a probability $p_{t}\colon E\rightarrow[0,1]$ defined on edges (for each round $t$), a budget $B$, and a time horizon $T$, a multi-round adaptive policy $\pi$ for IM-IC works as follows: at each round $t\in[T]$, let $S^{t}(\pi)$ be the set of seeds selected by $\pi$ in round $t$ and $A^{t}_{v}(\pi)$ be the set of nodes activated by $v\in S^{t}(\pi)$ (we will next omit the reference to $\pi$ when clear from the context). Note that both $S^{t}$ and $A^{t}_{v}$ are initially empty. Then, the policy at each round $t\in[T]$, if the total number of selected nodes is lower than $B$, works as follows:
@@
    It decides whether to choose another seed in the current round or to move to the next round (note that if all nodes have been activated in the current round, then the policy must necessarily choose to move to the next round).
@@
    In the first case, the policy selects a node $v$ that has not been yet activated in the current round, i.e., $v\notin A^{t}_{S^{t}}:=\bigcup_{u\in S^{t}}A^{t}_{u}$, observes the set $A^{t}_{v}$ of newly activated nodes, i.e., $A^{t}_{v}:=IC(G[V\setminus A^{t}_{S^{t}}],p,\{v\})$, adds $v$ to $S^{t}$, and repeat from step 1.
Finally, in the problem of Multi-Round Adaptive Influence Maximization under Independent Cascade diffusion model (MR-IM-IC) we are given a graph $G=(V,E)$, a time horizon $T$, a weight function $w_{t}\colon V\rightarrow\mathbb{R}_{+}$ defined on nodes for every $t\in[T]$, a probability $p_{t}\colon E\rightarrow[0,1]$ defined on edges for each round $t\in[T]$, and a budget $B$, and we are asked to select the multi-round adaptive policy $\pi^{*}$ that maximizes the expected sum of weights of activated nodes over $t$ rounds, i.e., $\pi^{*}=\arg\max_{\pi}\varsigma(\pi)$ where $\varsigma(\pi)=\mathbb{E}_{(p_{t})_{t\in[T]}}\left[\sum_{t\in[T]}\sum_{v\in A^{t}_{S^{t}}(\pi)}w_{t}(v)\right]$.
We stress that our definition allows that both nodes’ weights and edge probabilities may change from one round to the next: this can be useful to implement a variable discount factor (e.g., we care more about the number of people influenced in the last phases of the campaign) or to implement a dynamic environment in which the members of a population are known to appear and disappear over time, or the relevance of population members changes over time (e.g., in liquid democracy settings [7] at the beginning we are more interested in influencing a large portion of the population for manipulating the choice of delegates, while later in the campaign we are more concerned in influencing those members that have more chances to be chosen as delegates, in order to manipulate the outcome of their voting process).
@@
We start by showing how an instance of the MR-IM-IC problem can be mapped to an instance of SBSMm. Let $M=(T,B,V,E,(p_{t})_{t\in[T]},(w_{t})_{t\in[T]})$ be an instance of MR-IM-IC and let $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in[T]})$ denote the generic instance of SBSMm. The mapping works as follows:
  • $T$ and $B$ are exactly the same as in $M$ (hence, we use exactly the same name).
  • The set $V$ of items is the set of nodes in the social network.
@@
  • The set $H(v)$ of possible local states of $v$ corresponds to the possible sets of nodes influenced by an independent cascade starting from node $v$, i.e., $H(v)=\{X\subseteq V\mid x\text{ is reachable from }v\text{ in }G,\ \forall x\in X\}$; we observe that $|H(v)|$ is in general exponential in the input size, despite it can be succinctly represented in polynomial space.
  • For any global state $\eta_{t}\in H$ and $v\in V$, $\mathcal{P}_{t}(\eta_{t}(v))$ is the probability that an independent cascade influences exactly the $\eta_{t}(v)\in H(v)$ nodes when started from $v$. We can succinctly define $\mathcal{P}_{t}$ through the concept of live edges [32]. The live-edge graph $G^{L}_{t}=(V,E^{L})$ at round $t$ is a random subgraph of $G$ obtained by including each edge $(u,v)\in E$ in $E^{L}_{t}$ with probability $p_{t,(u,v)}$, independently of the random inclusion of the other edges; each edge $(u,v)\in G_{t}^{L}$ is called live as it can be used to spread the influence from $u$ to $v$, while each edge $(u,v)\notin G_{t}^{L}$ is called dead as it cannot do it. We observe that the probability of a given subgraph $G=(V,E^{\prime})$ to be the live-edge graph is $\mathbb{P}[G^{L}_{t}=G^{\prime}]=\prod_{(u,v)\in E^{\prime}}p_{t,(u,v)}\prod_{(u,v)\in E\setminus E^{\prime}}(1-p_{t,(u,v)})$. Then, $\mathcal{P}_{t}(\eta_{t}(v))$ can be equivalently defined as the sum of $\mathbb{P}[G^{L}_{t}=G^{\prime}]$, over all subgraphs $G^{\prime}=(V,E^{\prime})$ of $G$ such that the set of nodes reachable from $v$ in $G^{\prime}$ coincides with $\eta_{t}(v)$, i.e., $\mathcal{P}_{t}(\eta_{t}(v))$ is the probability that the set of nodes reachable from $v$ in $G_{t}^{L}$ is exactly $\eta_{t}(v)$.
  • The objective function $f_{t}$ is defined as $f_{t}(S,\eta_{t})=\sum_{v\in\bigcup_{u\in S}\eta_{t}(u)}w_{t}(v)$.
The mapping between a multi-round policy $\pi$ for IM-IC and a multi-round policy $\pi^{\prime}$ for SBMSm is determined as follows. The multi-round policy $\pi$, at a certain step of round $t$, either moves to the subsequent round, or adds a new seed $v\in V$ to the previously selected seeds set $S^{t}$ and observes the newly activated nodes. For the sake of simplicity, we will assume that $\pi$, after selecting an item $v$ as a seed, is even able to observe the entire set of nodes that $v$ would have activate in graph $G$ if no seed had been previously selected. This assumption is not restrictive, as knowing if a seed $v$ would have or not influence some already active nodes does not give useful information to the policy, and will be used below only to define the mapping with SBMSm. The corresponding multi-round policy $\pi^{\prime}$ of SBMSm, at a certain step of round $t$, either moves to the subsequent round, or adds a new seed $v\in V$ to the previously selected items $S^{t}$ and observes the local state $\eta_{t}(v)$. Considering that $\eta_{t}(v)$ represents the set of nodes influenced by a cascade starting from node $v$, the mapping between $\pi$ and $\pi^{\prime}$ easily follows.
@@
According to the mapping between $M$ and $I$ and that between the multi-round policies of such instances, it is immediate to check that $\varsigma(\pi)=\sigma(\pi^{\prime})$, and thus finding the policy maximizing the latter is equivalent to find a policy maximizing the former.
@@
In order to establish that $I$ is a valid instance for SBMSm we need also to prove that $f_{t}$ is polynomially computable and representable, it is polynomially bounded, and it satisfies monotonicity and adaptive submodularity. Polynomial representation and computation immediately follows from the definition above. As for polynomial boundedness, this property is satisfied if there exist constants $c,C>0$ such that $w_{t}(v)\leq C$ for every $v$ and every $t$, and there exist a round $t$ and a node $v$ such that $c\leq w_{t}(v)$; we observe that for standard unweighted influence maximization problems, where the objective function simply counts the number of influenced nodes, we can set $C=c=1$. Given the constants $c,C$ defined above, the values $\Lambda:=Cn$ and $\lambda:=c$ immediately satisfy the polynomial boundedness.
The properties of monotonicity and adaptive submodularity have been formally shown for the standard single-round adaptive influence maximization problem by [26], and clearly extend to the multi-round setting considered here. For the sake of completeness, we provide the main intuitions on how these properties can be checked. Monotonicity is immediately satisfied since the information diffusion does not decrease if the set of seeds expands. As for the adaptive submodularity, it is not hard to check that this condition is satisfied: intuitively, for each realization of the live-edge graph, as described above, the increment of $f_{t}$ due to the selection of a new seed $v$ is given by nodes reachable from $v$ in this live graph, but not reachable by previously selected seeds, and it is clearly impossible to expand this subset with more previously selected seeds.